Bob purchased 8 gallons of paint and 3 brushes for $152.50. The next day he purchased 6 gallons of paint and 2 brushes for $113.00 how much does each gallon of paint and each brush cost?
Answer by mananth(16946)

x= cost per gallon
y= cost per brush

8.00 x + 3.00 y = 152.50 .............1
Total value
6.00 x + 2.00 y = 113.00 .............2
Eliminate y
multiply (1)by -2.00
Multiply (2) by 3.00
-16.00 x -6.00 y = -305.00
18.00 x + 6.00 y = 339.00
Add the two equations
2.00 x = 34.00
/ 2.00
x = 17.00
plug value of x in (1)
8.00 x + 3.00 y = 152.50
136.00 + 3.00 y = 152.50
3.00 y = 152.50 -136.00
3.00 y = 16.50
y = 5.50
x= 17.00 cost per gallon
y= 5.50 cost per brush
m.ananth@hotmail.ca